# Deep Operator Networks for Bayesian Parameter Estimation in PDEs

### v1

@@ 3 Experiments
In this section, we train DeepBayONets on various algebraic and differential equations. In each case, the underlying equations have unknown parameters that are learned along with the forward dynamics. We compare the results with exact solutions and competing methodologies from the literature.
3.1 General Setup
The common configurations for all the experiments are as follows: each experiment utilizes a neural network architecture comprising three layers on both the branch and the trunk. The hyperbolic tangent ($\tanh$) activation function is employed universally. Training and optimization are conducted using the Adam optimizer. All implementations are performed in PyTorch, leveraging its advanced features for computational graphs and gradients. Specific details, such as the number of neurons per layer, epochs, batch size, and loss weights, are tuned for each experiment based on its requirements and complexity.
@@ 3.2 Forward Bayesian Regression Problem
This experiment demonstrates DeepBayONet’s capability to estimate aleatoric and epistemic uncertainty in a synthetic regression problem. The regression task is defined as follows:
@@
where $y_{i}$ are noisy observations with state-dependent inhomogeneous noise. The goal is to learn the function $f(x)$ while estimating two sources of uncertainty: Aleatoric uncertainty arising from state-dependent input noise in the data, and the epistemic uncertainty due to gaps in the data and testing outside the training range eq. 16.
The results are compared against various deep learning uncertainty quantification methods, including a Simple Neural Network with aleatoric uncertainty estimate (SNN), a Bayesian Neural Network (BNN), a Monte Carlo Dropout Network (MCDO), and a Deep Ensemble Network (DENN). Each architecture contains approximately 3,800–4,000 parameters to ensure a fair comparison. The data are split into a training set, an in-training-distribution (IDD) test set, and an out-of-training-distribution (OOD) test set.
The test problem and implementations are adopted from [36]. For a comprehensive review of uncertainty quantification methods in deep learning, refer to [37]. All networks are trained for 150 epochs with the Adam optimizer (batch size 16). Each model is run ten times with random initializations, and accuracies are averaged to mitigate random effects. Common hyperparameters are kept consistent across all models, while method-specific hyperparameters use default library values. To limit computational resource requirements, we have not done method-specific hyperparameter tuning.
@@ 3.4 One-dimensional unsteady heat equation
$w_{\text{IC}}=3,\quad w_{\text{data}}=6,\quad w_{\text{BC}}=1,\quad w_{\text{int}}=1,\quad w_{\text{param std}}=1.$
@@
Figure 4: Loss components during the training of the 1D Heat equation.
During training, several metrics were recorded to monitor the model’s performance and convergence, including the total loss and individual loss components: PDE (Interior) loss, boundary condition (BC) loss, initial condition (IC) loss, data loss, and the standard deviation loss for predicted parameters (STD loss). These metrics are shown in fig. 4, highlighting the stability and convergence of the model.
Figure 5 compares a sample of the posterior predicted solution $y(t,x)$ with the exact solution. The grainy appearance of the predicted solution reflects the uncertainty in the parameters $D$ and $\alpha$, which affects the pointwise forward solution. However, obtaining multiple samples of the output and computing statistical measures, such as the mean and mode, is computationally efficient once the model is trained.
@@ 3.5 Two-dimensional Reaction-diffusion equation
Figure 9: Loss history plot for the 2D Reaction-diffusion equation in eq. 19. We observe a decrease in all loss components except for the variance of the posterior parameter distribution, as intentionally designed.
The training history, as shown in fig. 9, demonstrates consistent reduction in the PDE (Interior) and data losses during early epochs, followed by a decrease in the boundary condition loss. The variance of the posterior parameter distribution remains relatively unchanged due to the KL divergence term in the loss function.
@@
Figure 12 shows the mean predicted solution and its error along with the predicted uncertainty for boundary-enforced training. In fig. 13, the corresponding plots for boundary-free training reveal reduced accuracy and elevated uncertainty, highlighting the ability of model to adapt to missing boundary conditions.
@@ 4 Conclusions
This study introduces Bayesian Deep Operator Networks (DeepBayONet) integrated with Physics-Informed Neural Networks (PINNs) to address parameter estimation and uncertainty quantification challenges in solving partial differential equations (PDEs). By employing a Bayesian framework with variational inference, the approach effectively captures both epistemic and aleatoric uncertainties, offering robust predictions in noisy and data-scarce environments. The method demonstrates versatility across tasks, including forward and inverse problems in one- and two-dimensional PDEs, with a relatively low-cost computational efficiency on par with maximum likelihood training as opposed to the traditional Bayesian neural network training.
The framework provides several advantages. It achieves comprehensive uncertainty quantification, enabling reliable predictions under complex scenarios, and maintains competitive accuracy. Additionally, the architecture integrates seamlessly with modern deep learning frameworks, allowing for scalability in addressing high-dimensional and nonlinear challenges.
However, the method has limitations. It requires careful tuning of hyperparameters, such as weights for loss components, which can introduce manual overhead. Although more efficient than many Bayesian alternatives, variational inference remains computationally intensive compared to deterministic methods.
Future work could address these challenges by exploring adaptive strategies for loss weighting to improve usability and generalizability. Further refinement of prior distributions to incorporate domain-specific knowledge could enhance parameter estimation and uncertainty quantification. We have focused on Gaussian priors in this work. Expanding the approach to complex parameter distributions and multiscale and multiphysics systems can broaden its applicability.
Acknowledgements
This research was supported by the CSML lab and the College of Engineering at California State University Long Beach. The work used Jetstream2 at Indiana University through allocation CIS230277 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.

### v2

@@ 3 Experiments
In this section, we train DeepBayONets on various algebraic and differential equations. In many cases, the underlying equations have unknown parameters that are learned along with the forward dynamics. We compare the results with exact solutions and competing methodologies from the literature when they are available.
3.1 General experiment setup
The common configurations used for all the experiments are as follows: each experiment utilizes a feedforward neural network architecture comprising three layers on both the branch and the trunk. The hyperbolic tangent ($\tanh$) activation function is employed universally. Training and optimization are conducted using the Adam optimizer. All implementations use the PyTorch package, leveraging its features for computational graphs and gradients. Specific details, such as the number of neurons per layer, epochs, batch size, and loss weights, are tuned for each experiment based on its requirements and complexity. These experiments are documented in the paper’s GitHub repository [^2] https://github.com/csml-beach/deep-bayesian-operator-nets for reference.
@@ 3.2 Forward Bayesian Regression Problem
This experiment demonstrates DeepBayONet’s capability to handle uncertainties in a synthetic regression problem. The regression task is defined as follows:
@@
where $y_{i}$ are noisy observations with state-dependent inhomogeneous noise. The goal is to learn the function $f(x)$ while estimating two sources of uncertainty: data(Aleatoric) uncertainty arising from state-dependent input noise in the data, and the model(epistemic) uncertainty due to the gap between the training data and inference points. Figure 2 shoes the dataset for this problem.
The results are compared against various deep learning uncertainty quantification methods, including a Simple Neural Network with aleatoric uncertainty estimate (SNN), a Bayesian Neural Network (BNN), a Monte Carlo Dropout Network (MCDO), and a Deep Ensemble Network (DENN). Each architecture contains approximately 3,800–4,000 parameters to ensure a fair comparison. The data are split into a training set, an in-training-distribution (IDD) test set, and an out-of-training-distribution (OOD) test set. The latter division is used to gauge the predicted uncertainty of the model outside the training range: Ideally, the model would become less confident the further away the inference point is from the training data. In fig. 2 the training range is denoted with vertical blue lines.
This test problem and the suite of uncertainty-aware deep neural networks are adopted from [38]. For a comprehensive review of uncertainty quantification methods in deep learning, refer to [39]. All networks are trained for 150 epochs with the Adam optimizer (batch size 16). Each model is run ten times with random initializations, and accuracies are averaged to mitigate random effects. Common hyperparameters are kept consistent across all models, while method-specific hyperparameters use default library values. To limit computational resource requirements, we have not performed method-specific hyperparameter tuning.
@@
Figure 2: (a) DeepBayONet without trunk (SNN). (b) Monte Carlo Dropout (MCDO). (c) Deep Bayesian Operator Network (DeepBayONet). Each method exhibits distinct uncertainty quantification behavior under the same experimental conditions. The two vertical blue lines in each plot denote the division between the IDD dataset (between the lines) and ODD dataset (everywhere else).
Figure 2 presents the uncertainty quantification for DeepBayONet without trunk (SNN), Monte Carlo Dropout Network, and regular DeepBayONet. SNN effectively predicts the uncertainty in the data within the training range accurately; however, it estimates relatively low uncertainty for data outside the training range. This behavior is attributed to the lack of any proper mechanism to quantify the uncertainty in the network parameters. On the other hand, MCDO predicts a higher uncertainty outside of the training range. This results from the activation perturbations induced by the dropout. Yet, MCDO exhibits overconfident prediction on inference points within the training range. Compared to figure 2a, which depicts DeepBayONet without the trunk, the regular DeepBayONet effectively captures meaningful uncertainty by perturbing the activation of the intermediate layer of the network. As a result, DeepBayONet accurately captures both sources of uncertainty for both data within and outside the training range as indicated in figure 2c.
@@ 3.4 One-dimensional unsteady heat equation
$w_{\text{IC}}=3,\quad w_{\text{Data}}=6,\quad w_{\text{BC}}=1,\quad w_{\text{Interior}}=1,\quad w_{\text{STD}}=1.$
@@
A basic grid search was employed to navigate the vast hyperparameter space and ensure the reproducibility of the experiments. Additionally, we implemented a learning rate scheduler that adaptively reduces the learning rate, further minimizing the model’s sensitivity to hyperparameters and improving its ability to navigate local minima. It’s worth noting that manually examining the magnitude of each component of the loss function eq. 17 between training epochs is also an option to adjust them proportionally.
During training, several metrics were recorded to monitor the model’s performance and convergence, including the total loss and individual loss components: PDE (Interior) loss, boundary condition (BC) loss, initial condition (IC) loss, data loss, and the standard deviation loss for predicted parameters (STD loss). These metrics are shown in fig. 5a, highlighting the stability and convergence of the model.
Figure 5: (a) Loss history plot for 1D Heat equation in eq. 21 eq. 21. (b) Loss history plot for the 2D Reaction-diffusion equation in eq. 22. (c) Loss history plot for 3D Helmholtz equation in eq. 23
Figure 6 compares a sample of the posterior predicted solution $y(t,x)$ with the exact solution. The grainy appearance of the predicted solution reflects the uncertainty in the parameters $D$ and $\alpha$, which affects the point-wise forward solution. However, obtaining multiple samples of the output and computing statistical measures, such as the mean and mode, is computationally efficient once the model is trained.
@@ 3.5 Two-dimensional Reaction-diffusion equation
Comparing the training loss history, as shown in fig. 5, we observe that the reaction-diffusion losses in fig. 5b have magnitudes of difference, while in fig. 5a the loss components are well-balanced. This is due to the fact that in fig. 5a the weights for different losses are about the same while the data loss and interior loss for the reaction-diffusion equation had to be scaled up dramatically for a successful training.
@@
Figure 12 shows the mean predicted solution and its error along with the predicted uncertainty for boundary-enforced training. In fig. 13, the corresponding plots for boundary-free training reveal reduced accuracy and elevated uncertainty, highlighting the ability of the model to adapt to missing boundary conditions.
@@
3.6 3D Eigenvalue problem in a unit ball
In this experiment, we use the three–dimensional Dirichlet–Laplacian 3D problem:
$\begin{cases}-\Delta u(\mathbf{x})=\lambda\,u(\mathbf{x}),&\|\mathbf{x}\|<1,\\[4.0pt] u(\mathbf{x})=0,&\|\mathbf{x}\|=1.\end{cases}$ (23)
In Cartesian coordinates the domain boundary $x^{2}+y^{2}+z^{2}=1$ cuts through a regular grid at oblique angles. Conventional grid-based solvers must either deform a Cartesian mesh to fit the curved boundary or use immersed-boundary tricks. Both options complicate stencil design. Furthermore, estimating the eigenvalue $\lambda$ as a parameter depends on the entire error in $u$, not just local residuals, which makes accurate estimation of both quantities challenging.
We solve this problem using DeepBayONet by approximating the eigenmode via the branch network and the eigenvalue using the trunk network. A numerical challenge in learning the eigenpairs is that the PDE and the boundary conditions are satisfied by $u\equiv 0$ for any $\lambda$, so additional constraints are required to force the network toward a non‐trivial eigenmode.
$\mathcal{L}=\underbrace{\mathbb{E}_{\Omega}\!\bigl{[}-\Delta u_{\theta}-\lambda_{\theta}u_{\theta}\bigr{]}^{2}}_{\text{PDE residual}}\;+\;\underbrace{\mathbb{E}_{\partial\Omega}\!\bigl{[}u_{\theta}^{2}\bigr{]}}_{\text{Dirichlet BC}}\;+\;\underbrace{\bigl{(}\mathbb{E}_{\Omega}[u_{\theta}^{2}]-1\bigr{)}^{2}}_{\text{normalization}}.$
The loss function for this problem combines three terms: (i) the PDE residual $(-\Delta u-\lambda u)^{2}$ evaluated at random interior points, (ii) the squared boundary condition on the sphere, and (iii) a simple normalization $(\mathbb{E}[u^{2}]-1)^{2}$ to guide the learned solution away from the trivial solution $u\equiv 0$. Automatic differentiation supplies exact gradients of all three terms, so the optimizer can update the network weights and the spectral parameter $\lambda$ in a single pass.
The eigen-functions of eq. 23 are available in closed form. They separate into spherical Bessel functions $j_{\ell}(k_{\ell s}r)$ and spherical harmonics $Y_{\ell}^{m}(\theta,\varphi)$. Our neural solver never sees this analytic structure. Throughout the experiments, the network receives only Cartesian coordinates $(x,y,z)$ and must satisfy the boundary condition $x^{2}+y^{2}+z^{2}=1$ implicitly through the loss. Working in Cartesian space, therefore, removes the coordinate system in which the PDE decouples, forces the optimizer to discover a non-trivial eigenfunction from scratch, and provides a more demanding test of the physics-informed loss. The first two eigenvalues are $\lambda_{01}=\pi^{2}\approx 9.87$ (radial mode) and $\lambda_{11}=4.49341^{2}\approx 20.19$ (dipole mode). During training, we monitor convergence toward $\lambda_{11}$.
Figure 14 presents a comprehensive visualization of the model’s performance on the 3D Helmholtz eigenvalue problem using both equatorial slices at fixed values and a meridional slice at $x=0$ In each slice, the model’s predicted field (left), the exact solution (center), and the point-wise absolute error (right) are shown. The polar plots for $z=0.0,0.25$ and $z=0.50$ illustrate that the model learns the correct dipole structure expected. The meridional slice, shown in the bottom row, further verifies the agreement between predicted and exact solutions in the full 3D domain. The errors are highly localized and remain within a small magnitude across the domain. To further assess the model’s accuracy along a 1D cut, Figure 15b shows the predicted and exact values of the solution along the central vertical axis $(x=y=0).$
Complementing this, the Figure 15a histogram of the learned eigenvalue parameter $\lambda$ confirms sharp convergence to the ground truth value $\lambda\approx 4.493$, validating that the network accurately identifies the underlying spectral structure.
Predicted
Exact
Error
$z=0.00$
$z=0.25$
$z=0.50$
$x=0$
Figure 14: Combined results for the 3D Helmholtz experiment.
(a) 3D Axis Plot
(b) 3D Histogram
Figure 15: Axis and Histogram plot for the 3D Helmholtz equation in eq. 23
@@ 4 Conclusions
This study introduces Bayesian Deep Operator Networks (DeepBayONet) integrated with Physics-Informed Neural Networks (PINNs) to address parameter estimation and uncertainty quantification challenges in solving partial differential equations (PDEs). By employing a Bayesian framework with variational inference, the approach effectively captures both data and model uncertainties, offering robust predictions in noisy and data-scarce environments. The method demonstrates versatility across tasks, including regression and PDE problems in one, two, and three dimensions.
The framework provides several advantages. It achieves comprehensive uncertainty quantification, enabling reliable predictions under complex scenarios, and maintains competitive accuracy. Additionally, the architecture integrates seamlessly with modern deep learning frameworks, allowing GPU-accelerated applications.
The method has limitations that may be addressed in future directions: Careful tuning of hyperparameter can be achieved algorithmically by adaptive strategies for loss weighting. Further refinements that can use non-Gaussian priors to incorporate domain-specific knowledge could enhance parameter estimation and uncertainty quantification. Finally, although DeepBayONet discovers the parameters of the PDE from the available data and physics, it does not learn a parametrized solution operator of the PDE. We leave that extension to future endeavors.
Acknowledgments
We thank the reviewers for their careful reading and thoughtful feedback, which helped us improve this work. This research was supported by the CSML lab and the College of Engineering at California State University Long Beach. The work used Jetstream2 at Indiana University through allocation CIS230277 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.